OpenStudy (anonymous):

Write the equation of the mirror line for the segment with endpoints (5, 4) and (-1, -1). Find the reflection of the point A(3, -2) through the mirror line of y = 1/2x − 1.

OpenStudy (polaris_s0i):

first find you're slope: \[ m = \frac{\Delta y}{\Delta x}\] \[ m = \frac{y_2 - y_1}{x_2 - x_1} \] \[ m = \frac{4 + 1}{5 + 1} = \frac{5}{6}\] then substitute a point into the equation: \[y = \frac{5}{6}x + b\] \[4 = \frac{5}{6}5 + b\] \[4 = \frac{25}{6} + b\] \[4 - \frac{25}{6} = b\] \[ b = -\frac{1}{6}\] so the equation of the mirror line is: \[y=\frac{5}{6}x - \frac{1}{6}\]

OpenStudy (anonymous):

wow thank you so much!o: do you know how to do the second one? Or can i do it in the same way you did the first one?

OpenStudy (polaris_s0i):

second one is different, one sec.

OpenStudy (polaris_s0i):

for reflection of (3,-2), you need to find the distance to the line: \[ y = \frac{1}{2}x - 1\] the "shortest" distance is on a line perpendicular to this one. perpendicular lines have a slope that is the negative reciprocal of the slope of the original so: \[ y = -2x + b \] plug in (3, -2) to find b: \[ -2 = -2(3) + b \] \[ -2 = -6 + b \] \[ 4 = b \] \[ y = -2x + 4 \] now you find the point where they intersect: \[ -2x + 4 = \frac{1}{2}x - 1 \] \[ -2x - \frac{1}{2}x = -1 - 4\] \[\frac{-4}{2}x - \frac{1}{2}x = -5\] \[\frac{-5}{2}x = -5\] \[x = 2\] the y coordinate is found by substituting x back into one of the lines (doesn't matter which one). \[y = -2*2 + 4\] \[y = -4 + 4\] \[y = 0 \] so the midpoint is (2, 0) this means we went one unit left in x to get to the midpoint, so we can go one more unit left to get to the reflection. So lets plug in 1 for x. \[y = -2(1) + 4\] \[y = 2\] so the reflection point should be (1, 2)

OpenStudy (polaris_s0i):

to check take calculate the distance between the midpoint and A: \[s^2 = (\Delta x)^2 + (\Delta y)^2 \] \[s^2 = (x_2 - x_1)^2 + (y_2 - y_1)^2\] \[s^2 = (3 - 2)^2 + (-2 - 0)^2\] \[s^2 = 1 + 4\] \[s = \sqrt{5}\] then calculate the distance from the midpoint to the reflected point: \[s^2 = (\Delta x)^2 + (\Delta y)^2 \] \[s^2 = (x_3 - x_2)^2 + (y_3 - y_2)^2\] \[s^2 = (2 - 1)^2 + (0 - 2)^2\] \[s^2 = 1 + 4\] \[s = \sqrt{5}\] they are the same, so we know we are correct!

OpenStudy (anonymous):

Thank you so much for these you're awesome! :3